Pretty sure it was bad diode.
Pretty sure it was bad diode.
So actually yes, I solved my problem. It wasn't the diodes. My problem was that I played a bit (a dangerous thing to do) with the menus and EEPROM contents of the controller while in "maintenance" mode and must have wrecked something. In the end I went through the RESET procedure of the laser head, then through all the calibration procedures again and that solved it.
In your case, I would:
1/ go through an LBO temp optimization procedure. The automated way never worked for me, so I would put the laser controller in maintenance mode, then in current control mode, start the laser at an intermediate current level, then scan a large range of temperatures of the LBO while monitoring the laser power output. The Coherent VERDI software is very useful for this
2/ go through the FAP modules temp optimization procedures. The automated way never worked for me either, so I would put the laser controller in maintenance mode, then in current control mode, start the laser at an intermediate current level, then scan a large range of temperatures of the diodes while monitoring the laser power output. The Coherent VERDI software is very useful for this, same as above
3/ (if 1/ not successfull) ensure the diodes are working properly (/!\ hat requires tooling and IR eye protection /!\). Basically you open up the controller housing, unplug a fiber from a FAP module, place an optical power meter at the exit of a diode (which can handle up to 50W), put on the safety goggles, place the laser in maintenance mode, place the diodes in current mode, then slowly ramp up the current on the diodes to see if you get the intended optical power out of the FAP modules
4/ if 2/ shows that the diodes are fine but still no output, I would go through a head reset procedure. Start however by exporting/recording all laser parameters with the Coherent VERDI software.
All steps above at your own risk, it normally requires a skilled, trained and tooled engineer![]()
Thank you. Unfortunately, I am a novice. I did try the LBO temperature optimization through command prompt, but that did not help. Wanted to run FAP-I optimization but not sure how to access the diode optimization menu/service login on GUI. Also forgot to add in my earlier post, I am using this Verdi V18 to pump a chameleon, my pump is around 12-14W, but the chameleon head stays at 0mW and is giving me a lasing fault (error 52; low output power). All which started to happen just recently when laser wavelength was changed suddenly while it it was trying to modelock.
Sorry, I don't know the Chameleon system, only the Verdi.
Hello
I encountered a problem with Verdi 10W.
I turned off the power supply with the LBO cooling-down procedure
for about a week
After I turned it on, I observed
1. An error code # 4 LBO temperature fault. The set point for LBO temperature read 38400 . The LBO temperature was raised up to ~180 oC.
2. After restarting the PS, I observed that Vanadate set point temperature read 553402322211286548000.0 (this is not a joke; this is precisely what the screen read )
There was also an issue with the battery, Error code 30, but I fixed it by replacing the old PS1212 )
3. Unlocking the service menu did not help much; the LBO setpoint reacted on the knob; of course, it would take hours to scroll it down. But, the set point for Vanadate did not respond to the knob rotating at all.
4. After pressing the motherboard against the slot, plugging and unplugging the service menu jumper, and turning on/off the PS, I finally got “reasonable” values for LBO and Vanadate setpoints, which could be adjusted.
5. Unfortunately, the service report with temperature settings was lost.
Coherent representative suggested (do they keep service reports stored? )
To use the following settings
LBO 148.8
Vanadate 30
Etalon 36
Diode 1 =23.8
Diode 2 =17
Later, I found the problem with LBO temperature oscillation in the 144.6 - 152.0 oC range without any tendency for stabilization for hours. In the LBO servo menu, I found PID settings and changed
the proportional gain from 10 to 250,
the Integ. Gain form 60 to 67
the Diff. gain remains the same 3.
After those adjustments, I could get a stable LBO temperature, with a stability of +-.2 C. (is there any self-adjustment procedure? )
6. The highest value of “set power” I can set is 2.2 W (is it the max that can be used as a set power? Before the problem occurred, the setpoint was 6 W) I tried to run the system with the settings I got but failed. After turning the key on and opening the shutter, the system stayed in the laser-seeking mode for an hour. With no progress
7. A coherent representative suggested that I switch to the current control mode and try to optimize the settings manually. So, with the initial settings, the maximum power I got was 0.3 W .
Playing with the setting, I could get the desired 9 W power.
Now I got
LBO 149.3 C
Diode 1 and Diode 2 both have ~ 26.1 C
Vanadate temperature 30
However, there is an issue with these settings.
I could stabilize power at ~ 9 W level, but starting the system the next day requires me to optimize settings again, mainly Etalon temperature, and raise it. So now Etalon’s temperature is 42 C, but a few days ago, it was 36 C.
I can work only in the Current control mode.
I also noticed that changing the current does not change the power smoothly.
The power reads 0.05 W (I think it is an error ) from 0 A up to ~ 22, then rises very slowly up to 0.2~ 0.3 W with the current increasing up to 25.2 A; after that, it jumps up to ~ 6 W. and I can get ~ 8.7 with current 26.6A After that I play with Etalon setting to reach 9W
I do not mind operating the system in its current mode, but I do not like that I need to optimize the temperature settings every time I run it. The current-power dependency (sharp jump around 25.2A) also does not look healthy to me.
Any thoughts on what can be checked or optimized? I would be grateful for any ideas. Maybe someone can help with access to the service manual?
Cheers,
IK
Sounds like there are issues with the control electronics in the power supply. Id get some deoxit and spray all of the connectors. But the weird vanadate reading is odd. It might be the memory was corrupted.
Hi
Thank you for the quick response!
By the connectors do you mean the contact pads going to the motherboard slot?
Regarding the memory
I also believe that memory was corrupted.
If I turn off/on the system some settings, etalon and Diodes temperatures, reset to their "default"(?) values. But it keeps the new LBO set point.
is there a separate chip which can be replaced to fix memory?
Thanks
IK